OpenStudy (anonymous):

a car accelerates uniformly from rest to a speed of 22.8 km/h in 8.1s -find the distance in travels during this time (answer i units of m)

OpenStudy (anonymous):

22.8 km/h*(1000m/1km)*(1h/3600sec)=6.333m/sec

OpenStudy (anonymous):

v(average)=1/2 (v(final)+v(initial) distance=v(average)*t

OpenStudy (anonymous):

Given initial velocity u=o m/s final velocity v= 22.8 km/h =6.33 m/s Given acceleration is uniform & time as 8.1s => acceleration a = (v-u)/t = 6.33/8.1 = 0.781 m/s^2 Distance travelled during this time = 1/2(at^2) = 1/2(0.781* 8.1^2) = 25.62m